the sum of two numbers is 38 when 8 is added to twice one of the numbers the result is five times the other numbers.find the two numbers.thank you
Answer by macston(5194)

.
F=first number; S=second number
.
F+S=38
S=38-F . Use this to substitute for S below.
.
2F+8=5S . Substitute for S from above.
2F+8=5(38-F)
2F+8=190-5F
7F=182
F=26
ANSWER 1: The first number is 26.
.
S=38-F=38-26=12
ANSWER 2: The second number is 12.
.
CHECK:
2F+8=5S
2(26)+8=5(12)
52+8=60
60=60